OpenStudy (11study11):

Help with 3 questions Need to finish is 30 min.!!!!

OpenStudy (11study11):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 8(1.05)n Part A: When the scientist concluded his study, the height of the plant was approximately 11.26 cm. What is a reasonable domain to plot the growth function? Part B: What does the y-intercept of the graph of the function f(n) represent? Part C: What is the average rate of change of the function f(n) from n = 2 to n = 6, and what does it represent?

OpenStudy (11study11):

@jim_thompson5910 please help!!!

jimthompson5910 (jim_thompson5910):

is that second 'n' an exponent? or no?

OpenStudy (11study11):

exponent

OpenStudy (11study11):

Part A: Domain will be talking about the values of n. Now when we solve for f(n) = 11.26 we get 7.005, So a reasonable range to plot the graph would be from n = 0 to 8 Part B: The y-intercept is the point where n = 0. But n represents the number of days, i.e. time elapsed. So at n = 0 the point represents the initial starting point. Part C: At n = 2, f(2) = 8*(1.05)^2 = 8.82 At n = 6, f(6) = 8*(1.05)^6 = 10.72 The average rate of change is (change in height)/(change in n) The change in height is 10.72 - 8.82 = 1.9 The change in n is 6 - 2 = 4 So the average rate of change is 1.9/4 = 0.475 This is what I have so far is it correct?

jimthompson5910 (jim_thompson5910):

Everything is correct. Nice job.

OpenStudy (11study11):

The price of products may increase due to inflation and decrease due to depreciation. Derek is studying the change in the price of two products, A and B, over time. The price f(x), in dollars, of product A after x years is represented by the function below: f(x) = 12500(0.82)x Part A: Is the price of product A increasing or decreasing and by what percentage per year? Justify your answer. Part B: The table below shows the price f(t), in dollars, of product B after t years: t (number of years) 1 2 3 4 f(t) (price in dollars) 5600 3136 1756.16 983.45 Which product recorded a greater percentage change in price over the previous year? Justify your answer. (5 points)

OpenStudy (11study11):

Part A: This function is decreasing because 0.63, the number inside of the parenthesis, is less than 1. 1 - 0.63 = 0.37. So it is decreasing by 0.37. Part B what I have so far.

jimthompson5910 (jim_thompson5910):

where are you getting 0.63? I see 0.82

OpenStudy (11study11):

ops

OpenStudy (11study11):

Part A: This function is decreasing because 0.82, the number inside of the parenthesis, is less than 1. 1 - 0.82 = 0.18. So it is decreasing by 0.18.

jimthompson5910 (jim_thompson5910):

decreasing by 18%

OpenStudy (11study11):

ok

OpenStudy (11study11):

Part B?

OpenStudy (11study11):

@jim_thompson5910

OpenStudy (11study11):

@erinkb99

OpenStudy (wwhitlock):

Math like this gives me a headache. I should go now.